OpenStudy (anonymous):

Is y = 2 + 4x a linear equation?

jimthompson5910 (jim_thompson5910):

y = 2 + 4x y = 4x + 2 it is in the form y = mx+b where m = 4 and b = 2

jimthompson5910 (jim_thompson5910):

so it is definitely a linear equation